1 0
Add a comment Improve this question Transcribed image text
Answer #1

Fixed Cost = Total Cost when output is 0

Q

TC

MC

FC

VC

AVC

ATC

0

60

60

0

1

160

100

60

100

100

160

2

220

60

60

160

80

110

3

270

50

60

210

70

90

4

340

70

60

280

70

85

5

450

110

60

390

78

90

6

630

180

60

570

95

105
Answer #2
Output(Q) total cost (TC)

MC= delta TC/delta Q

FC = TC-VC    VC AVC=VC/Q ATC=TC/Q
0 60 --- $60 0 -- ----
1 160 $100 $60 $100 $100 $160
2 220 $60 $60 $160 $80 $110
3 270 $50 $60 $210 $70 $90
4 340 $70 $60 $280 $70 $85
5 450 $110 $60 $390 $78 $90
6 630 $180 $60 $570 $95 $105